there is a narrow rectangular plot, reserved for a school, in mahuli village. the length and breadthof the plot are in the ratio 11:4. at the rate Rs 100 per meter it will cost the village panchayat Rs 75000 to fence the plot. What are the diemensions of the plot?
Answer by mananth(16946)

there is a narrow rectangular plot, reserved for a school, in mahuli village. the length and breadth of the plot are in the ratio 11:4.
let the common multiple be x
the length will be 11x
breadth = 4x
Perimeter = 2(L+B)
=2(11x+4x)
= 2*15x
=30x
at the rate Rs 100 per meter
30x * 100 = 3000x
3000x= 75000
3x=75
x=25
11x = 11*25 = 275m
4x=4*25 = 100 m
275m & 100m are the dimensions of the plot